A small sphere of charge q1 = +8.00 x 10-5 C is at a distance r = 2.67 m from another small sphere of charge q2 = -6.00 x 10-5 C (a) What is the magnitude of the attractive electric force between these two charged spheres? (b) How many excess protons are there in the small sphere with charge q1 ? (c) How many excess electrons are there in the small sphere with charge q2 ?


1
Expert's answer
2021-02-22T10:22:57-0500

Answer

(a)

Attractive force

"F=k\frac{q_1q_2}{r^2}"


"F=(9\times10^9)\frac{(8.0\times10^{-5})(6.0\times10^{-5})}{2.67^2}"


"F=6N"

(B)

Number of proton on small sphere

"q_1=n_1e\\n_1=\frac{q_1}{e}=\frac{8.00\times10^{-5}}{1.6\times10^{-19}}=5\times10^{14}"


(C)

Excess of electron on sphere

"q_2=n_2e\\n_1=\frac{q_1}{e}=\frac{6.00\times10^{-5}}{1.6\times10^{-19}}=3.75\times10^{14}"